# Dynamic LLM-Agent Network: An LLM-agent Collaboration Framework with Agent Team Optimization

### v1

@@ 3.2 Construction of DyLAN
To further enhance efficiency, we introduce an early-stopping mechanism. Inspired by the Byzantine Consensus theory (Castro & Liskov, 1999), at least $13p+$ 1 agents are needed to tolerate $p$ faulty agents in a single round of communication. Following the theory, the inference process will be terminated when over 2/3 of agents in a single layer have a consistent answer. In practice, the inference process will also be terminated when the maximum time step is reached. Note that none of the consistency measures used in prior work (Wang et al., 2023a; Aggarwal et al., 2023) applies to multi-round multi-agent interaction since their theories are assumed to execute a single LLM instance multiple times. In our case, we apply the early-stopping mechanism to each layer of DyLAN.
In short, DyLAN gives the answer in a forward passing manner with an efficient dynamic architecture for agent collaborations (Algorithm 1). Other implementation details are in Appendix C.1.
3.3 Agent Team Optimization
To automatically find the optimal team of agents among candidates for specific domains and based on their actual responses, we set up an algorithm in which the agent team optimization process could be formulated as a function $f_{\mathrm{Optim}}:S\rightarrow S^{\prime}$, where $S^{\prime}=({\mathbb{A}}^{\prime},\mathbb{E}^{\prime})$ and ${\mathbb{A}}^{\prime}$ denotes a smaller set of agents with better performance. With a pool of agent candidates, we implement $f_{\mathrm{Optim}}$ in a three-step procedure, as shown in Figure 1: (1) Propagation: Take all candidates into a collaboration, ask each node to rate the solutions to the task query from its predecessors (in the forward pass); (2) Aggregation: Each node aggregates the ratings it has received from its successors towards itself (via another backward pass) to quantify its own contribution independently at different time steps. (3) Selection: During the last step, we sum up the scores for the same agent over all time steps to derive an importance score for each agent, and extract the top-$k$ agents that are most contributory according to these scores.
@@ 3.3 Agent Team Optimization
Specifically, at the step of Propagation, for each agent, we ask it to rate the responses from all its predecessors. Formally, the $i$-th agent at the $t$-th time step takes its prompt $p_{i}$, the input query $q$, and all previous responses ${\mathbb{R}}_{t-1}$ from their predecessors, and further map them via a scoring function $f^{(s)}_{t,i}(\cdot,\cdot,\cdot)$ to produce the rating scores. Here, we use $w_{t-1,j,i}$ to refer to the rating score on $a_{t-1,j}$ from $a_{t,i}$, and $[w_{t-1,1,i},w_{t-1,2,i},...,w_{t-1,n,i}]=f^{(s)}_{t,i}(p_{i},q,{\mathbb{R}}_{t-1})$. After propagation, the contribution of node $a_{t,i}$ is the sum of its successors’ contribution multiplied by their peers’ ratings on the agent’s response. Formally, the aggregation process is described as:
${\bm{I}}_{t-1,j}=\sum_{(a_{t-1,j},a_{t,i})\in\mathbb{E}}{\bm{I}}_{t,i}\cdot w_{t-1,j,i},$ (1)
@@
where ${\bm{I}}_{t,i}$ denotes the contribution of $a_{t,i}$. In the final Selection step, Agent Importance Score ${\bm{I}}_{i}$ for the $i$-th agent is defined as ${\bm{I}}_{i}=\sum_{t=1}^{T}{\bm{I}}_{t,i}$. In practice, we initialize the contributions in the final layer first, and step backward to perform Aggregation layer by layer (Algorithm 2). The definition guarantees that the agent importance scores add up to 1 in each layer, which benefits fair comparison. Other details, such as initializing contributions in the final layer, are presented in Appendix C.2.
@@
Table 2: Accuracy (%) on MATH dataset. CoT refers to Chain-of-Thought prompting (Wei et al., 2022) and examples are from original dataset. The number in parentheses indicates the performance difference relative to a single execution. The median of three trials is reported when non-zero temperature is used.
@@
Table 3: Accuracy (%) on MMLU dataset. “Other” stands for subjects like business, health, and misc in MMLU. The median of three trials is reported when non-zero temperature is used.
Method Humanities Social STEM Other Overall #API
Science Calls
@@
DyLAN (Ours) 62.1 79.1 69.7 75.5 70.5 (+4.1) 4.39
4 Experiments
4.1 Setup
To verify the effectiveness and efficiency of DyLAN, we conduct extensive experiments on three representative tasks including arithmetic reasoning, general reasoning, and code generation.
Arithmetic Reasoning We leverage MATH (Hendrycks et al., 2021b) as the evaluation dataset, which consists of 7 subareas and contains 5,000 questions in the test set. We choose LLM Debate (Du et al., 2023), LLM-Blender (Jiang et al., 2023), and the single execution on LLM as baselines. We also compared DyLAN to the previous state-of-the-art method PHP (Zheng et al., 2023). To draw a fair comparison, we categorize systems by prompting strategies, including normal CoT prompts (Wei et al., 2022) provided in the original dataset and Complex CoT provided by PHP. Since the task is sensitive to hyper-parameters, we have tuned and used each system’s best temperature. Preliminary experiments show that collaborating agents of different roles did not introduce significant improvement, therefore we adopt agents of the same role on this dataset for all systems.
General Reasoning For the general reasoning task, we use the MMLU dataset (Hendrycks et al., 2021a), which contains four aspects of a vast amount of problems in 57 subjects. We down-sample 1/5 of the problems in the test set because of its huge quantity. We use the same baselines as arithmetic reasoning. To align with previous works and prevent degeneration of LLM Debate, we set the hyper-parameter temperature to 1 for all systems. For each subject, 7 agents with different roles are used as the candidates when evaluating DyLAN, and 4 of them are selected to form the optimized agent team based on Agent Importance Score. The roles of candidates match the categories of MMLU, including “Mathematician” and “Programmer” for STEM, “Lawyer” and “Historian” for Humanities, “Economist” and “Psychologist” in Social Science, and “Doctor” for clinical questions in the “Other” category. We extract the prompts from an open-source codebase and details are provided in Appendix E.
Code Generation We use the HumanEval benchmark in the code generation task, with 164 human-labeled function-level completion codes and unit tests (Chen et al., 2021). We found that LLM Debate and LLM-Blender could not feasibly adapt to this task in their implementation, thus they are excluded them from the baselines. As a result, we leverage two strong methods CodeT (Chen et al., 2023a) and Reflexion (Shinn et al., 2023) along with the single execution as the baselines. For agent team optimization in DyLAN, 12 candidates are given and the team size is set to 8. Additionally, we require 4 of the 8 agents should be the role of code writer and the other 4 are judges providing code reviews. And two of the judges are individually augmented with a code interpreter and a syntax checker. Please refer to Appendix C.1 for more details.
@@ 4.2 Main Results
Table 4: Experimental results on HumanEval dataset. We indicate the foundation model of methods except for GPT-35-turbo. The median of three trials is reported when non-zero temperature is used.
Method Pass@1 #API
Calls
Single Execution 73.2 (+0.0) 1.00
CodeT 65.8 (-7.4) 20.00
CodeT (Codex) 74.8 (+1.6) 20.00
Reflexion 68.3 (-4.9) 4.05
DyLAN (Ours) 82.9 (+9.7) 16.85
Single Execution (GPT-4) 88.4 (+15.2) 1.00
Reflexion (GPT-4) 91.4 (+18.2) 7.32
DyLAN (Ours, GPT-4) 92.1 (+18.9) 15.94
In Table 2, Table 3, and Table 4, we report the classification accuracy on MATH and MMLU, and Pass@1 on HumanEval, respectively. Additionally, we recorded the average times of calling LLMs of each method on each query as #API calls. This metric serves as a proxy for the computational cost and the efficiency of architectures for LLM-agent collaborations, which cannot be determined from the number of tokens since the number varies greatly depending on the query.
DyLAN improves overall performance in different tasks. Table 2 shows that DyLAN has significantly higher performance than baselines on arithmetic reasoning, demonstrating the advantage of the dynamic architecture. Relative to a single execution, it gains a +4.1 improvement with CoT prompts and +3.5 with Complex CoT prompts, demonstrating robustness against different prompting strategies. In Table 3 and Table 4, DyLAN also significantly improves the performance by +4.1 and +9.7 on general reasoning and code generation tasks, respectively, relative to the single execution. We argue part of the improvement can be attributed to the dynamic multi-path architecture, which allows different opinions to be delivered simultaneously. In contrast, for methods in sequential architecture like PHP (Zheng et al., 2023), incorrect intermediate answers might easily influence the final output due to only one reasoning path. It is the same for Reflexion (Shinn et al., 2023) in code generation, in which false tests may mislead the self-debugging process from our observations. In our case, any feedback from predecessors could be rated by the nodes at any time step, making it easier to identify false intermediate feedback explicitly.
DyLAN has a reasonable computational cost. From Table 2, we find DyLAN realizes an 10.2% improvement to LLM Debate in terms of accuracy, with 10.6% lower #API calls (L3 vs. L4), suggesting it is a better trade-off between efficiency and effectiveness. Similar trends can be observed in Table 3. DyLAN has a better overall accuracy with only 36.6% API calls of LLM Debate (4.39 vs. 12.00). Moreover, we see that DyLAN dynamically adjusts the cost based on the difficulty of a query. Given that questions in MMLU are less challenging than in MATH, DyLAN has 2.76 fewer #API calls on the query from MMLU compared to one from MATH. This makes sense as generally, the more complex the task, the harder it will be to come to a consensus.
Table 5: The optimized composition of roles and performance improvement in different subjects of MMLU dataset when optimizing DyLAN of 7 agents to 4 agents.
@@
DyLAN benefits from agent team optimization. Additionally, we found that an optimized team of agents could enhance DyLAN. For different subjects in MMLU, agent compositions are adjusted correspondingly to improve up to 25.0% in accuracy, as shown in Table 5. After agent team optimization, we find a smaller set of agents are selected from the given large pool of candidates for specific subjects, resulting in significant performance improvement, such as “Mathematician”, “Programmer”, and “Economist” for college mathematics, and “Doctor” and “Psychologist” for clinical knowledge. Moreover, before performing agent team optimization, i.e., all the 12 agents are used, the performance of DyLAN on HumanEval is 76.2, which means agent team optimization introduces an improvement of +6.7 points (76.2 $\rightarrow$ 82.9), suggesting that Agent Importance Scores can effectively capture and reflect the actual abilities of agents to particular task domains.
4.3 Ablation Studies
[Uncaptioned image][Uncaptioned image] Figure 2: Experimental results of optimizing agent team with small subsets of MMLU (left) and HumanEval (right). X-axis denotes the proportion to the original datasets. We also randomly selected the same amount of agents as a baseline.
Data Efficiency of Agent Team Optimization We further demonstrate the data efficiency of agent team optimization by performing it based on different amount of data. The experiments are conducted on the MMLU and HumanEval datasets. We sample the subsets with the proportions of 0.01 and 0.1 of the original dataset. Agent team optimization is performed on the subsets and tested the optimized team of agents on the whole dataset. The settings are identical to experiments in Section 4.1. As shown in Figure 2, with the optimized team of agents on 0.1 of the original dataset, DyLAN has demonstrated similar performance compared to optimizing on the whole dataset, with only 0.2 loss of overall accuracy on MMLU and 0.6 loss of Pass@1 on HumanEval. Especially, as code review has significant impact on the performance on code generation, it is critical to select a proper set of judges to provide code review. We can observe that even with only 0.01 of the original dataset, DyLAN could obtain a significant improvement of +3.7 over random selection on HumanEval, indicating its effectiveness on selecting proper team of agents.
Impact of Optimized Agent Team Size As shown in Figure 3, DyLAN with an optimized team of 3 agents can outperform both the same architecture with 7 agents and LLM Debate with 4 agents, suggesting the effectiveness of our proposed agent optimization. The efficiency is also significantly improved by 52.9% and 67.8%, respectively. These results indicate that an optimized selection of agents would be able to collaborate on tasks better and reach the consensus much faster and more accurately than many agents thrown together at random.
Figure 3: Impact of optimized agent team size. 2$\sim$4 agents are selected from 7 candidate agents based on Agent Importance Score. Accuracy (left) and #API calls (right) on MMLU are reported.
Table 6: Ablation study on the early-stopping mechanism (es) and the inference-time agent selection component (its). #API denotes the average number of API calls.
Method MATH MMLU HumanEval
Acc. #API Acc. #API Pass@1 #API
DyLAN 35.7 7.15 70.5 4.39 82.9 16.85
w/o es 35.0 13.00 70.1 13.00 80.5 19.00
w/o its 33.8 8.20 69.9 7.05 76.2 17.98
@@ 4.3 Ablation Studies
Impact of Early-Stopping and Inference-Time Agent Selection As shown in Table 6, early-stopping mechanism boosts efficiency to a great extent by minimizing #API calls by 45.0%, 66.2%, and 11.3% on MATH, MMLU, and HumanEval respectively, while providing slight performance improvement. Inference-time agent selection, however, is critical to enhance the correctness of the final answer. We conjecture it is because low-performing agents are filtered at specific time steps.
@@
Stability of DyLAN with Different Backbone Models There is also a notable difference in code generation tasks when the backbone model changes (Table 4). Reflexion and CodeT’s performances are heavily related to the backbone model (L4 vs. L5 and L6 vs. L9). Instead, DyLAN shows a steady, consistent high performance (L7 vs. L10) under different backbone models with almost the same amount of API calls.

### v2

@@ 3.3.2 Agent Team Reformation
To further enhance efficiency, we introduce an early-stopping mechanism. Inspired by the Byzantine Consensus theory (Castro & Liskov, 1999), at least $13p+$ 1 agents are needed to tolerate $p$ faulty agents in a single round of communication. Following the theory, the inference process will be terminated when over 2/3 of agents in a single layer have a consistent answer. In practice, the inference process will also be terminated when the maximum time step is reached. Note that none of the consistency measures used in prior work (Wang et al., 2023b; Aggarwal et al., 2023; Yin et al., 2023) applies to multi-round multi-agent interaction since their theories are assumed to execute a single LLM instance multiple times or expect all agents to reach the same answer.
3.4 Team Optimization
The goal of team optimization is to select a subset of agents from candidates based on their contributions evaluated from a primary trial, such that the new team solves the task query more effectively and efficiently. Formally, given a task query $q$, a set of agents $\mathcal{A}$, a trial is performed based on the algorithm proposed in Section 3.3.2 resulting in a T-FNN $\mathcal{G}_{q}^{\mathcal{A}}$. And team optimization is formulated as
$\hat{\mathcal{A}}=f_{\mathrm{Optim}}(\mathcal{A},\mathcal{G}_{q}^{\mathcal{A}},q),\text{ where }\hat{\mathcal{A}}\subset\mathcal{A}.$ (9)
$f_{\mathrm{Optim}}$ is implemented as in a three-step procedure of agent selection (Figure 2 right):
@@ 3.4 Team Optimization
(1) Propagation: Each node rates the solutions to the task query from its predecessors, which is a forward message passing process. Formally, given a node $v_{t,j}$ and an edge $(v_{t-1,i},v_{t,j})$, the message $m_{v_{t-1,i}}$ sent from $v_{t-1,i}$ to $v_{t,j}$ is defined as the response to the task query $q$ from the agent $a_{i}$ at the previous time step. The aggregator function $f_{\mathrm{mp}}(\cdot,v_{t,j})$ is implemented as a scoring function $f^{(s)}_{t,j}(\cdot,\cdot,\cdot)$, which maps the prompt $p_{j}$, the input query $q$, and all the messages $\mathcal{M}$ to the rating scores. Here, we use $w_{t-1,i,j}$ to refer to the rating score on $v_{t-1,i}$ from $v_{t,j}$, and
$[w_{t-1,1,j},w_{t-1,2,j},...,w_{t-1,N,j}]=f^{(s)}_{t,j}\left(p_{j},q,\mathcal{M}\right).$ (10)
(2) Aggregation: Each node aggregates the ratings it has received from its successors towards itself to quantify its own contribution independently at different time steps. The contribution of node $v_{t-1,i}$ is the sum of its successors’ contribution multiplied by their peers’ ratings on the agent’s response. Aggregation is a backward message passing process. Formally, given a node $v_{t-1,i}$ and an edge $(v_{t-1,i},v_{t,j})$, the message $m_{v_{t,j}}$ sent from $v_{t,j}$ to $v_{t-1,i}$ is defined as $w_{t-1,i,j}$. And the aggregator function $f_{\mathrm{mp}}$ is defined as a weighted sum function:
$\bm{I}_{t-1,i}=\sum_{(v_{t-1,i},v_{t,j})\in E_{t-1,t}}\bm{I}_{t,j}\cdot w_{t-1,i,j},$ (11)
where $\bm{I}_{t,i}$ denotes the contribution of $a_{t,i}$.
(3) Selection: During the last step, we sum up the scores for the same agent over all time steps to derive an importance score for each agent, and extract the top-$k$ agents that are most contributory according to these scores to form the optimized agent team. Formally, the Agent Importance Score $\bm{I}_{i}$ for agent $a_{i}$ is defined as
$\bm{I}_{i}=\sum_{t=1}^{T}\bm{I}_{t,i}.$ (12)
@@
In practice, we initialize the contributions in the final layer first, and step backward to perform Aggregation layer by layer (Algorithm 2). The definition guarantees that the agent importance scores add up to 1 in each layer, which benefits fair comparison. Other details, such as initializing contributions in the final layer, are presented in Section B.2.
Method Pass@1 #API Calls
Single Execution 73.2 (+0.0) 1.00
CodeT 65.8 (-7.4) 20.00
CodeT (Codex) 74.8 (+1.6) 20.00
Reflexion 68.3 (-4.9) 4.05
LATS 81.1 (+7.9) 48.00
CAMEL 69.5 (-4.1) 12.03
AgentVerse 75.0 (+1.8) 22.50
DyLAN (Ours) 82.9 (+9.7) 16.85
Method Reward Success #API
Rate Calls
Direct Execution 50.6 (+0.0) 28.0 14.52
ReAct 53.8 (+3.2) 30.0 8.40
ReAct-SC 58.0 (+7.4) 36.0 25.75
Reflexion (trial=4) 62.0 (+11.4) 40.0 25.40
LATS 64.5 (+13.9) 38.0 $>$ 400
BOLAA 66.0 (+15.4) 40.0 32.40
DyLAN (Ours) 68.3 (+17.7) 42.0 24.85
@@
Table 2: Experimental results on the CG task (left) and results on the DM task (right). The number in parentheses indicates the difference relative to the single execution or direct execution. We indicate the foundation model of methods except for GPT-35-turbo. The median of three trials is reported when non-zero temperature is used.
@@
Table 3: Accuracy (%) on the AR task. The number in parentheses indicates the performance difference relative to a single execution.
4 Experiments
4.1 Setup
Code Generation (CG) We use the HumanEval benchmark, with 164 human-labeled function-level completion codes and unit tests (Chen et al., 2021). Unit tests are used to validate the correctness of generated codes. We leverage two strong baselines CodeT (Chen et al., 2023a) and Reflexion (Shinn et al., 2023) along with the single execution. For multi-agent baselines, we re-implement CAMEL (Li et al., 2023) and AgentVerse (Chen et al., 2024) under their original configurations for fair comparisons.
Decision Making (DM) We evaluate our methods in the WebShop environment, selecting 50 environments in its test set (Chen et al., 2021) following the setting of LATS (Zhou et al., 2023). WebShop requires to find the item given an instruction of the customer. It provides “reward” as an intrinsic metric for item-instruction relevance, and “success” is marked when the reward is 1.0. Besides ReAct (Yao et al., 2023) and Reflexion, we re-ran a multi-agent method BOLAA (Liu et al., 2023), and a single-agent method LATS as strong baselines.
General Reasoning (GR) For the general reasoning task, we use the MMLU dataset (Hendrycks et al., 2021a), which contains four categories of a vast amount of problems in 57 subjects. We down-sample 1/5 of the problems in the test set because of its huge quantity. We choose LLM Debate (Du et al., 2023), LLM-Blender (Jiang et al., 2023), and the single execution on LLM as baselines.
Arithmetic Reasoning (AR) We leverage the test set of MATH (Hendrycks et al., 2021b) for evaluation, which consists of 7 subareas and 5,000 questions in total. To draw a fair comparison and verify the robustness, we categorize methods by different prompting strategies and select strong baselines accordingly. Preliminary experiments show that collaborating agents in different domains (e.g., algebra and geometry experts) does not make significant improvement, therefore we adopt agents with same prompts for all methods.
DyLAN Setup In Table 5, we elaborate the setup of DyLAN. To keep in line with baseline methods, we only equipped DyLAN with code interpreters as tools in the CG task. It is worth noting that agent selection is performed for each subject in the GR task, for each web page in the DM task, and directly for CG task in the team optimization stage. Due to space limitations, please refer to Section B.1 for details.
Method Hum- Social STEM Other Overall #API
anities Science Calls
@@ 4.1 Setup
DyLAN 62.1 79.1 69.7 75.5 70.5 (+4.1) 4.39
Table 4: Accuracy (%) on the GR task. “Other” stands for subjects like business, health, and misc in the MMLU dataset. We report the median of three runs for experiments.
Task #Agents Tool Performance #API Calls
Usage Improvement
CG $12\rightarrow 8$ 76.2 $\rightarrow$ 82.9 23.04 $\rightarrow$ 16.85
DM $8\rightarrow 4$$\times$ 53.0 $\rightarrow$ 68.3 32.03 $\rightarrow$ 24.85
GR $7\rightarrow 4$$\times$ 69.5 $\rightarrow$ 70.5 08.30 $\rightarrow$ 04.39
AR $4$$\times$ - -
Table 5: Demonstration of experiment settings, including the number of agents and the performance throughout team optimization. We report reward for the DM task.
@@ 4.2 Main Results
In Table 2, Table 3, and Table 5, we report the results on each dataset respectively. The number of API calls serves as a proxy for the efficiency of communication structures for agents, which cannot be clearly determined from token consumption which varies greatly depending on tasks and prompting strategies. Since “Task solving” after “Team Optimization” is essential for testing and deployment, we mainly report the cost from the second stage. The difference can be seen in Table 5, and further discussions in Section C.1.
DyLAN improves overall performance on different tasks with a reasonable computational cost. From Table 3, we find DyLAN realizes an 10.2% improvement to LLM Debate in terms of accuracy, with 10.6% lower #API calls (L3 vs. L4), suggesting it is a better trade-off between efficiency and effectiveness. Similar trends can be observed as DyLAN has a better performance with only 36.6% API calls of LLM Debate (L5 vs. L4 in Table 5), and 35.1% of LATS on CG and $<$6% on DM (L8 vs. L5 (left), L7 vs. L5 (right) in Table 2). We argue it can be attributed to the feed-forward structure and early-stopping mechanism, which allows different solutions to be delivered simultaneously and confirmed rapidly. In contrast, for methods in sequential architecture like PHP and Reflexion (Table 2), incorrect intermediates might easily influence the final output due to the single thread of reasoning or code generation and review. Also, ReAct on DM tasks exhibits similar failures due to misoperations in the middle. In contrast, Reflexion and LATS access the environment at certain states for multiple times to for reflection, limiting generalizabilities. In our case, any feedback from predecessors could be rated by successor nodes, making it easier to rectify potential invalid actions. Moreover, we see that DyLAN dynamically adjust the cost based on the difficulty of tasks. For instance, most questions in the MMLU dataset are less challenging than MATH, DyLAN has 2.76 fewer API calls on the query from the former. However, compared to other tasks, DyLAN introduces relatively lower improvements in AR tasks, which might be due to the high knowledge dependency of the MATH dateset.
DyLAN benefits from the team optimization. Moreover, we found that a dynamic team of task-oriented agents could enhance DyLAN. For different subjects in GR tasks, agent compositions are adjusted correspondingly to improve up to 25.0% in accuracy, as shown in Table 8. As denoted in Table 5, a dynamically selected team of agents could result in significant performance improvement, especially for DM tasks, where agents might have great interference from others. The overall performance can be significantly improved (up to 6.7%) with lower computational costs on each tasks after agent selection. Moreover, it also suggests that Agent Importance Scores can effectively capture and reflect the actual contributions of agents on a wide range of tasks. We further verify this claim in Section C.6.
Figure 3: Impact of optimized agent team size. 2$\sim$4 agents are selected from 7 candidate agents based on Agent Importance Score. Accuracy (left) and #API calls (right) on the GR task are visualized.
Method AR GR
Acc. #API Acc. #API
DyLAN 35.7 7.15 70.5 4.39
w/o es 35.0 13.00 70.1 13.00
w/o atr 33.8 8.20 69.9 7.05
Method CG DM
Pass@1 #API Reward #API
DyLAN 82.9 16.85 68.3 24.85
w/o es 80.5 19.00 67.5 54.25
w/o atr 76.2 17.98 66.0 48.90
Table 6: Impact off the early-stopping mechanism (es) and agent team reformation (atr).
4.3 Ablation Studies
Impact of Optimized Agent Team Size Fewer proper agents in a team could have better performance. As shown in Figure 3, DyLAN with an optimized team of 3 agents can outperform both DyLAN before team optimization and LLM Debate with 4 agents, suggesting the effectiveness of our proposed agent selection. The efficiency is also significantly improved by 52.9% and 67.8%, respectively. Probably because the imbalance of agents’ expertise and opinions interfere with each other before optimization, especially on GR, where few candidates are relevant to a subject.
@@ 4.3 Ablation Studies
Table 7: The optimized composition of agents and performance improvement on different subjects in the GR task.
#Code #Code Pass@1 #API
Writers Reviewers Calls
6 6 76.2 23.04
4 4 82.9 16.85
4 3 81.1 14.64
4 2 77.4 12.50
3 3 78.0 11.73
3 2 75.6 09.60
Table 8: Different compositions of agents on the CG task of an optimized team of agents. Agent teams are optimized by the Agent Importance Score from the first line.
Robustness of Agent Importance Score The Agent Importance Score is robust over the imbalance of agent roles. On GR tasks, the candidates are imbalanced in terms of expertise. For most queries, there are less than 2 candidates that are related according to their role prompts. In Table 8, we found agent selection is capable for selecting related agents, e.g., “Mathematician” for “college mathematics”, that matches human priors. However, if candidates are all vastly different from the task domain, e.g., “public relation”, where the improvement is less significant. We also tested DyLAN on CG tasks with different amount of code writers and code reviewers after the “Team Optimization” stage. It is worth noting that a single run of “Team Optimization” could provides reusable Agent Importance Scores for multiple trials of agent selection. In Table 8, we exhibit the results under imbalanced teams of agents. We verified the imbalance of code writers and reviewers after optimization won’t cause great performance drops. Though, reviewers affect the performance slightly greater than writers (L2,3 vs. L4,5), indicating the necessity of the amount of reviewers for code refinement.
@@
Impact of Early-Stopping and Agent Team Reformation As shown in Table 6, early-stopping mechanism boosts efficiency to a great extent by minimizing #API calls by 45.0%, 66.2%, 11.3%, and 54.2% on AR, GR, CG, and DM tasks respectively, while providing slight performance improvement. Agent team reformation, however, is critical to enhance the correctness of the final answer. We conjecture it is because agents are filtered for temporary mistakes in LLMs, such as hallucinations, etc. Additionally, answer comparison is more challenging for open-ended tasks like CG or DM tasks. We use the BLEU score with a 0.9 threshold for consistency checks. This makes early stopping less effective since agents may generate codes in different formats, leading to fewer opportunities to stop early.
@@
Stability of DyLAN with Different Backbone Models There is also a notable difference in CG tasks when the backbone model changes (Table 2). Reflexion and CodeT’s performances are heavily related to the backbone model (L4 vs. L5 and L6 vs. L9). Instead, DyLAN shows a steady, consistent high performance (L7 vs. L10) under different backbone models with almost the same amount of API calls.